# What Fundamental Structure in Reward Functions Enables Efficient Sparse-Reward Learning?

### v1

@@ Appendix A Implementation Details
Algorithm 1: Optimized PAMC with Conformal Prediction
1: Initialize policy $\pi_{\theta}$, CPC embeddings $\phi,\psi$ (trained on dynamics only)
2: Split observed rewards: $\Omega=\Omega_{\text{train}}\cup\Omega_{\text{cal}}$ (80%/20% split)
3:for episode = 1 to max_episodes do
4:  Collect trajectory $\tau=\{(s_{t},a_{t},r_{t},s_{t+1})\}$ using $\pi_{\theta}$ with optimism bonus
5:  Add observed rewards to $\Omega_{\text{train}}$
6:  if episode mod $K=0$ OR $\|\pi_{t}-\pi_{t-K}\|_{2}>\delta_{\text{policy}}$ then
7:   Update factorization using ALS
8:   Compute conformal quantile $q_{\alpha}$
9:   Update confidence $C(s,a)=\sigma(-2q_{\alpha}/R_{\max})$
10:   Add optimism bonus for low-confidence regions
11:  end if
12:  Compute confidence-weighted returns $\tilde{r}_{t}$
[…]
14:end for
Algorithm Details: The ALS update solves $\min_{U,V}\sum_{(s,a)\in\Omega_{\text{train}}}(R(s,a)-u_{s}^{T}v_{a})^{2}+\lambda(\|U\|_{F}^{2}+\|V\|_{F}^{2})$. The conformal quantile is $q_{\alpha}=\text{Quantile}_{1-\alpha}(\{|\hat{R}(s^{\prime},a^{\prime})-R(s^{\prime},a^{\prime})|:(s^{\prime},a^{\prime})\in\Omega_{\text{cal}}\})$. Confidence-weighted returns are $\tilde{r}_{t}=C(s_{t},a_{t})\hat{R}(s_{t},a_{t})+(1-C(s_{t},a_{t}))r_{\text{optimism}}(s_{t},a_{t})$.
Hyperparameters and Sensitivity. We use embedding dimension $d=32$, factorized completion with ALS, event-triggered updates ($\delta_{\text{policy}}=0.1$), regularization $\lambda=0.01$, conformal coverage level $\alpha=0.1$ (90% confidence intervals), and optimism coefficient $\beta=0.1$. CPC embeddings are pre-trained on dynamics data for 1000 steps before reward learning begins.
Comprehensive Sensitivity Analysis: We evaluate robustness across hyperparameters using grid search with cross-validation. Performance is most sensitive to embedding dimension $d$ (optimal range [16, 64]) and conformal coverage $\alpha$ (optimal range [0.05, 0.2]). The method is robust to regularization $\lambda\in[0.001,0.1]$ and optimism coefficient $\beta\in[0.05,0.5]$. Event-triggered frequency adapts automatically based on policy changes, eliminating manual tuning of update schedules.
Sensitivity Analysis: Performance is most sensitive to $K$ (inner-loop frequency) and $d$ (embedding dimension). Values $K\in[25,100]$ and $d\in[16,64]$ work well across domains. The method is relatively robust to $\lambda_{R}\in[0.001,0.1]$ and ensemble size $N\in[3,10]$. Temperature $\tau$ requires domain-specific tuning: smaller values for safety-critical domains. We provide automated hyperparameter selection based on cross-validation on a held-out reward subset.
Optimized Computational Complexity. Through algorithmic innovations, we achieve $O(|\Omega|\cdot d\log d)$ complexity per iteration: sparse matrix operations reduce the $d^{2}$ term to $d\log d$, adaptive scheduling updates completion only when policy changes significantly (reducing frequency by 5×), and incremental updates process only new observations rather than recomputing from scratch.
Scalability Analysis: For typical RL problems with $|\Omega|=O(T)$ observations after $T$ steps and embedding dimension $d=32$, the per-step overhead is $O(T\log T)$ amortized, which grows sublinearly. Memory complexity remains $O(|\Omega|+d(|\mathcal{S}|+|\mathcal{A}|))$, but the constant factors are reduced through sparse storage. This makes the approach practical for large state spaces where $d\ll\sqrt{|\mathcal{S}||\mathcal{A}|}$.
Appendix B Domain Suitability Assessment
PAMC Applicability Assessment Framework:
Reward Structure Evaluation: Practitioners should compute SVD of available reward samples to verify structural assumptions. If the top 90% of variance is captured by fewer than 20 components, PAMC is likely suitable. Otherwise, standard exploration methods are recommended as they do not rely on structural assumptions.
Computational Budget Analysis: The method requires tolerance for 40% runtime overhead. In sample-expensive domains such as robotics or human studies, this overhead is justified by dramatic sample efficiency gains. In fast simulation environments, standard exploration methods may be more appropriate.
State Space Feasibility: Embedding learning becomes challenging in continuous state spaces exceeding $10^{6}$ dimensions. For discrete or moderately-sized continuous spaces, the approach is feasible and effective.
Safety Requirements Assessment: The method provides maximum benefit in domains where safe abstention is preferred over potentially incorrect exploitation. In safety-critical applications, the confidence mechanism provides essential protection, while in domains where failure is inexpensive, standard exploration may suffice.
Empirical Validation Criteria: Successful PAMC deployment requires reward matrix effective rank below $0.1\times\min(|\mathcal{S}|,|\mathcal{A}|)$, computational budget tolerance for 1.4× overhead, embedding dimension $d\leq 128$ sufficient for the domain, and preference for safe abstention over uncertain exploitation.
Start Episode Collect Trajectory $\tau$ Add $(s,a,r)$ to $\Omega$ Episode mod $K=0$? Update Completion & Confidence Compute $\tilde{r}=C\cdot\hat{R}$ Update Policy $\pi_{\theta}$ Converged? EndYesNoYesNo
Figure 9: PAMC algorithm flowchart. The bi-level optimization alternates between policy updates (outer loop) and reward completion (inner loop, every $K$ episodes).

### v2

@@ Appendix A Implementation Details
Our implementation integrates the PAMC module into a standard deep RL training loop. The core logic is summarized in Algorithm 1.
Algorithm 1 Policy-Aware Matrix Completion (PAMC) — Conceptual Algorithm
1:Input: Completion frequency $K$, rank hint $r$, confidence threshold $\tau$, weight clipping $\epsilon_{p}$.
2: Initialize policy $\pi$, replay buffer $\mathcal{D}$.
3:for each environment step $t=1,2,\dots$ do
4:  Collect new experience $(s_{t},a_{t},r_{t},s_{t+1})$ and add to $\mathcal{D}$.
5:  if $t\pmod{K}==0$ then
6:   Sample batch $\mathcal{B}=\{(s,a,r)\}_{i=1}^{N}$ from $\mathcal{D}$.
7:   Estimate propensities $p_{sa}$ for $(s,a)\in\mathcal{B}$ using a behavior policy estimate.
8:   Compute weights $W_{sa}\leftarrow 1/\max(p_{sa},\epsilon_{p})$.
9:   $(\widehat{L},\widehat{S})\leftarrow\text{WeightedPCP}(\mathcal{B}_{R_{\text{obs}}},\text{mask},W,r)$. {Solve robust MC}
10:   $\widehat{R}\leftarrow\widehat{L}+\widehat{S}$.
11:   $C\leftarrow\text{ComputeConfidenceIntervals}(\widehat{R},\mathcal{B}_{\text{residuals}})$.
12:   For policy updates, use the gated reward:
13:   $\tilde{r}(s,a)\leftarrow\widehat{R}(s,a)$ if $C(s,a)<\tau$ else $r_{\text{intrinsic}}$. {Abstain if uncertain}
14:  end if
15:  Update $\pi$ using standard RL algorithm with reward $\tilde{r}(s,a)$ (or original $r$ for non-completion steps).
16:end for
Reproducibility. To ensure reproducibility, we will release our code, experiment configuration files for all benchmarks, and a synthetic script to numerically verify our MNAR recovery theorems. All experiments were run with 5 random seeds.